Expected Portfolio Return Beta 22 0.8 A Market 173 1.0 D) Expected Portfolio Return Beta 30.28 1.8 A Market 198 1.0If the simple CAPM is valid and all portfolios are priced correctly, which of the situations below is possible? Consider each
Answer #1

Option (D) is possible situation.

If CAPM is valid then expected return of stock and required return of stock under CAPM should be same.

In case of option-D,

Stock A expected return = 22%

And,

Market return = 19%

Risk free rate = 5%

Stock A beta = 1.8

Using CAPM equation we compute the required return of Stock A:

T = risk free rate + beta* (Market return - risk free rate)

r = 0.05 + 1.8*(0.19 -0.05)

T = 0.05 + 1.8*(0.14)

r = 0.05 + 0.252

T= 0.302

T = 30.2%

required return under CAPM = Expected return, thus CAPM is valid here. In other cases, CAPM is not valid.

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.